OpenStudy (anonymous):

ıntegrate(xcosx+5^2+3e^3x) plz help me

OpenStudy (anonymous):

\[\int\limits_{}^{}(x \cos(x)+e^{3x}+25) \ dx \]Is this what you have? @horlaryemzy

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

Well do it term by term:\[\int\limits_{}^{}(x \cos(x)+e^{3x}+25) \ dx =\int\limits_{}^{}x \cos(x) \ dx+ \int\limits_{}^{} e^{3x} \ dx+ \int\limits_{}^{}25 \ dx \]Each of these can be done but we'll need to use integration by parts/tabular integration to integrate xcos(x). The other two integrals are straight forward.

OpenStudy (anonymous):

thankz @genıus10

OpenStudy (anonymous):

np

OpenStudy (anonymous):

ı almost got ıt @horlaryemzy

OpenStudy (anonymous):

oryt @jeddydoncuzy